what is the maximum number of concurrent normals to an ellipse?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans: 4
Hello student,
Please find answer to your question below
Equation of ellipse:
\frac{x^{2}}{a^{2}} + \frac{y^{2}}{b^{2}} = 1
Equation of normal:
\frac{ax}{cos\theta } - \frac{by}{sin\theta } = a^{2}-b^{2}
Let the normal passes through point (x1, y1)
\frac{ax_{1}}{cos\theta } - \frac{by_{1}}{sin\theta } = a^{2}-b^{2}\frac{ax_{1}(1+tan^{2}\frac{\theta }{2})}{1-tan^{2}\frac{\theta }{2} } - \frac{by_{1}(1+tan^{2}\frac{\theta }{2})}{2tan\frac{\theta }{2} } = a^{2}-b^{2}
This polynomial equation in theta have four real & distinct roots.
So maximun four normal is possible.